The speed of light in ice is 2.3 X 10^8 m/s. What is the
refractive index of ice?
1
Expert's answer
2020-10-25T18:28:10-0400

By the definition of the refractive index we have:


"n=\dfrac{c}{v},"

here, "n" is the refractive index of ice, "c=3\cdot 10^8\ \dfrac{m}{s}" is the speed of light in a vacuum, "v=2.3\cdot 10^8\ \dfrac{m}{s}" is the speed of light in ice.

Then, we get:


"n=\dfrac{c}{v}=\dfrac{3\cdot 10^8\ \dfrac{m}{s}}{2.3\cdot 10^8\ \dfrac{m}{s}}=1.304"

Answer:

"n=1.304"